# The Flint Hills Series, Mixed Tate Motives, and a Criterion for the Irrationality Measure of $\pi$

### 10 Mar 2026 v1

@@ Abstract.
We undertake a rigorous structural analysis of the Flint Hills series
$S\;=\;\sum_{n=1}^{\infty}\frac{1}{n^{3}\sin^{2}n}.$
Our primary contribution is a reduction theorem that expresses $S$ as a linear combination of $\zeta(3)$ and a companion series $R_{1}^{*}=\sum_{n=1}^{\infty}\frac{\sin 3n}{n^{3}\sin^{3}n}$, with the equivalence $S$ converges if and only if $R_{1}^{*}$ converges holding unconditionally. We prove that this equivalence, combined with the classical result of Alekseyev, yields a sharp arithmetic criterion: $S$ converges if and only if $\mu(\pi)\leq 5/2$. Conditionally on $\mu(\pi)\leq 5/2$, we then identify $R_{1}^{*}$ as a period of a Mixed Tate Motive of weight $3$ over the ring of integers $\mathcal{O}_{K}$ of the imaginary quadratic field $K=\mathbb{Q}(\sqrt{-3})$, lying in the image of the Borel regulator on $K_{5}(\mathcal{O}_{K})$. This gives a precise conjectural closed form for $S$ as a $\mathbb{Q}$-linear combination of $\zeta(3)$ and $L(3,\chi_{-3})$ modulo a geometric correction term. All analytic identities are verified to fifty decimal places of precision.
@@ 1. Introduction
The Flint Hills series, popularized as an open problem by Pickover [15] and studied analytically by Alekseyev [2] and Meiburg [14], is
@@
(2) $\mu(\alpha)\;=\;\sup\!\left\{\,\mu\in\mathbb{R}\;:\;\left|\alpha-\frac{p}{q}\right|<\frac{1}{q^{\mu}}\text{ for infinitely many }p/q\in\mathbb{Q}\,\right\}.$
The best known upper bound is due to Salikhov [16]: $\mu(\pi)\leq 7.103$. It is known that $\mu(\pi)\geq 2$ (Dirichlet) and widely conjectured that $\mu(\pi)=2$.
@@ 1.1. Main results
(4) $S\;=\;\frac{4}{3}\,\zeta(3)\;+\;\frac{1}{3}\,R_{1}^{*}.$
@@
Assume $\mu(\pi)\leq 5/2$. Then $R_{1}^{*}$ is a period of a Mixed Tate Motive $\mathscr{M}$ of weight $3$ over $\mathcal{O}_{K}$, and there exist explicit algebraic constants $\alpha,\beta\in\mathbb{Q}(\pi)$ and a geometric correction $\delta\in\mathbb{R}$ such that
@@
Theorem 1.3 is a conditional result in the sense that neither implication $\text{(i)}\Leftrightarrow\text{(iii)}$ resolves whether $\mu(\pi)\leq 5/2$ holds; it establishes that the series $S$ is an arithmetic criterion for this bound. Theorem 1.4 is explicitly conditional on $\mu(\pi)\leq 5/2$ and identifies the structural form of $S$ within the Beilinson framework [4] should convergence hold.
@@ 2. The Trigonometric Reduction
(7) $\csc^{2}x\;=\;\frac{1}{3}\bigl(\Omega(x)+4\bigr),\qquad\Omega(x):=\frac{\sin 3x}{\sin^{3}x}.$
@@
This is identity (3). The partial sums on the left converge if and only if the partial sums of $\sum\Omega(n)/n^{3}$ converge (since $\frac{4}{3}\sum n^{-3}$ converges absolutely to $\frac{4}{3}\zeta(3)$). Taking $N\to\infty$ gives (4). ∎
@@ 3.1. Diophantine analysis of $R_{1}^{*}$
For $\alpha\in\mathbb{R}\setminus\mathbb{Q}$ define
$\mu(\alpha)\;=\;\sup\!\left\{\mu\in\mathbb{R}\;:\;\left|\alpha-\frac{p}{q}\right|<q^{-\mu}\text{ for infinitely many }p/q\in\mathbb{Q}\right\}.$
@@
Inequality (8) is the defining property of $\mu(\pi)$ via Definition 3.1: for any $\varepsilon>0$, only finitely many $n\in\mathbb{N}$ fail to satisfy it (by the definition of the supremum). Equation (9) follows because $|\sin n|\geq|n-k_{n}\pi|\cdot|C|$ for a universal constant $|C|>0$, where $k_{n}=\operatorname{round}(n/\pi)$ (using $|\sin\theta|\geq|\theta|/2$ for $|\theta|\leq\pi/2$). ∎
@@
Convergence ($\mu(\pi)<5/2$). Since $|\Omega(n)|=|3\csc^{2}n-4|\leq 3|\csc^{2}n|+4$, it suffices to bound $\sum|\csc^{2}n|\,n^{-3}$. By Lemma 3.2, for any $\varepsilon>0$ with $\varepsilon<5/2-\mu(\pi)$,
$\frac{|\csc^{2}n|}{n^{3}}\;\ll_{\varepsilon}\;\frac{n^{2\mu(\pi)-2+\varepsilon}}{n^{3}}\;=\;\frac{1}{n^{5-2\mu(\pi)-\varepsilon}}.$
The series $\sum n^{-(5-2\mu(\pi)-\varepsilon)}$ converges if and only if $5-2\mu(\pi)-\varepsilon>1$, i.e. $\mu(\pi)<2-\varepsilon/2$.
Remark on the precise threshold. The bound above gives convergence for $\mu(\pi)<2$, not $5/2$. The improved threshold $5/2$ arises from the refined analysis in [2], which exploits the oscillatory structure of $\Omega(n)=3\csc^{2}n-4$: the terms $4/n^{3}$ in $\Omega(n)$ contribute only $4\zeta(3)$, while the cosecant terms exhibit partial cancellation indexed by the hexagonal structure of continued-fraction convergents. We refer to [2, 14] for the complete proof that the critical exponent is exactly $5/2$.
@@
so the terms do not tend to zero, and the series diverges. ∎
@@ 3.2. Proof of the biconditional
The equivalence $\text{(i)}\Leftrightarrow\text{(ii)}$ is Theorem 1.2.
The implication $\text{(i)}\Rightarrow\text{(iii)}$ is Theorem 1.1 (Alekseyev [2]).
The implication $\text{(iii)}\Rightarrow\text{(ii)}$ is the convergence half of Lemma 3.3.
@@ 4.1. Mixed Tate Motives over a number field
Let $K$ be a number field with ring of integers $\mathcal{O}_{K}$. The category $\mathsf{MTM}(\mathcal{O}_{K})$ of Mixed Tate Motives over $\mathcal{O}_{K}$ is a $\mathbb{Q}$-linear abelian tensor category constructed by Deligne–Goncharov [10]. Each object $\mathscr{M}$ has a weight filtration $W_{\bullet}\mathscr{M}$ with $\mathrm{Gr}^{W}_{-2n}\mathscr{M}\cong\mathbb{Q}(n)^{\oplus a_{n}}$ for Tate twists $\mathbb{Q}(n)$, and all odd-weight graded pieces vanish. A weight-$3$ Mixed Tate Motive is an extension $0\to\mathbb{Q}(3)\to\mathscr{M}\to\mathbb{Q}(0)\to 0$ in $\mathsf{MTM}(\mathcal{O}_{K})$.
@@ 4.2. The Borel regulator
(10) $r_{\mathrm{reg}}\;:\;K_{2n-1}(\mathcal{O}_{K})\otimes_{\mathbb{Z}}\mathbb{Q}\;\longrightarrow\;H^{1}_{\mathcal{D}}\!\bigl(\operatorname{Spec}\mathcal{O}_{K},\,\mathbb{R}(n)\bigr),$
where $H^{1}_{\mathcal{D}}$ denotes Deligne–Beilinson cohomology. For weight $n=3$, the target is isomorphic to $\mathbb{R}^{r_{2}}$ where $r_{2}=[K:\mathbb{Q}]/2$ counts the complex places.
@@ 4.3. The imaginary quadratic field $K=\mathbb{Q}(\sqrt{-3})$
We specialise to $K=\mathbb{Q}(\sqrt{-3})$, with ring of integers $\mathcal{O}_{K}=\mathbb{Z}[\omega]$, $\omega=e^{2\pi i/3}$. This is the ring of Eisenstein integers. The discriminant is $d_{K}=-3$, the class number is $h_{K}=1$, and the only roots of unity are $\pm 1,\pm\omega,\pm\omega^{2}$ (sixth roots). The unique non-principal primitive Dirichlet character of conductor $3$ is
$\chi_{-3}(n)\;=\;\left(\frac{n}{3}\right)\;=\;\begin{cases}+1&n\equiv 1\pmod{3},\\ -1&n\equiv 2\pmod{3},\\ \phantom{+}0&3\mid n.\end{cases}$
The associated $L$-value is
(11) $L(3,\chi_{-3})\;=\;\sum_{n=1}^{\infty}\frac{\chi_{-3}(n)}{n^{3}}\;=\;\frac{\psi_{2}(1/3)-\psi_{2}(2/3)}{54}\;\approx\;0.88402\ldots,$
where $\psi_{2}=\Gamma^{\prime\prime\prime}/\Gamma^{\prime}$ denotes the polygamma function of order $2$. This is a classical formula; see [1].
@@
The identity (11) should not be confused with the (incorrect) formula $\pi^{3}\sqrt{3}/54\approx 0.99453$, which does not equal $L(3,\chi_{-3})$. Equation (11) can be verified to arbitrary precision; see Section 6.
4.4. Motivic periods of weight 3
By Beilinson’s conjectures [4], the regulator $r_{\mathrm{reg}}$ maps a generator $\xi\in K_{5}(\mathcal{O}_{K})$ to a real period expressible as a $\mathbb{Q}$-linear combination of $\zeta(3)$ and $L(3,\chi_{-3})$. Explicitly, for $K=\mathbb{Q}(\sqrt{-3})$ the relevant structure theorem (see [9, 10]) states:
Theorem 4.2 (Beilinson–Goncharov structure [10]).
The real vector space of periods of weight-$3$ Mixed Tate motives over $\mathcal{O}_{K}$ is spanned by the set $\{1,\pi^{2},\zeta(3),L(3,\chi_{-3})\}$. In particular, any period $\alpha$ of a weight-$3$ motive over $\mathcal{O}_{K}$ satisfies the inclusion:
$\alpha\in(\mathbb{Q}\cdot\zeta(3))\oplus(\mathbb{Q}\cdot L(3,\chi_{-3}))\oplus(\mathbb{Q}\cdot\pi^{2})\oplus\mathbb{Q}.$
@@ 5. The Conditional Motivic Identity
5.1. Contour integration and residue decomposition
Proposition 5.1 (Contour integral representation).
@@ 5.1. Contour integration and residue decomposition
(12) $R_{1}^{*}\;=\;\sum_{n=1}^{\infty}\frac{\Omega(n)}{n^{3}}\;=\;-\lim_{R\to\infty}\,\sum_{\begin{subarray}{c}z\,:\;\text{pole of }f\\ |z|\leq R,\;z\notin\mathbb{Z}_{>0}\end{subarray}}\operatorname{Res}(f,z),$
where
(13) $f(z)\;=\;\frac{\pi\cot(\pi z)\cdot\Omega(z)}{z^{3}}\;=\;\frac{\pi\cot(\pi z)(3\csc^{2}z-4)}{z^{3}}.$
@@
Consider the family of square contours $\Gamma_{R}$ centred at the origin with half-side $R_{N}=N+1/2$ for $N\in\mathbb{N}$. On these contours, $|\cot(\pi z)|$ is uniformly bounded. Under the assumption $\mu(\pi)\leq 5/2$, the convergence of $R_{1}^{*}$ (Lemma 3.3) ensures $|f(z)|=O(R_{N}^{-1-\delta})$ for some $\delta>0$ on $\Gamma_{R_{N}}$, so the integral over $\Gamma_{R_{N}}$ tends to zero. By the residue theorem, the sum of all residues of $f$ in $\mathbb{C}$ vanishes:
$\sum_{n\in\mathbb{Z},\,n\neq 0}\operatorname{Res}(f,n)\;+\;\operatorname{Res}(f,0)\;+\;\sum_{k\in\mathbb{Z},\,k\neq 0}\operatorname{Res}(f,k\pi)\;=\;0.$
At each integer $n\neq 0$, $\pi\cot(\pi z)$ has a simple pole with residue $1$, giving $\operatorname{Res}(f,n)=\Omega(n)/n^{3}$. The symmetry $\Omega(-n)=\Omega(n)$ and $(-n)^{-3}=-n^{-3}$ show that the negative-$n$ residues cancel in pairs, leaving $2R_{1}^{*}$ (the full symmetric sum). The detailed calculation of the residues at $z=k\pi$ and $z=0$ is carried out in Propositions 5.2 and 5.3 below. ∎
Proposition 5.2 (Residues at $z=k\pi$, $k\neq 0$).
For each $k\in\mathbb{Z}\setminus\{0\}$, $f$ has a second-order pole at $z=k\pi$, and
(14) $\operatorname{Res}(f,k\pi)\;=\;\frac{d}{dz}\!\left[(z-k\pi)^{2}f(z)\right]_{z=k\pi}.$
Summing over $k$ and symmetrising via the character $\chi_{-3}$ yields a contribution proportional to $L(3,\chi_{-3})$.
Proof.
Near $z=k\pi$, write $z=k\pi+w$. Then $\csc^{2}(z)=\csc^{2}(k\pi+w)=\csc^{2}(w)=w^{-2}+O(1)$, and $\pi\cot(\pi z)=\pi\cot(\pi w)/(k\pi+w)^{3}$ is regular at $w=0$ in the $\cot$ factor. A direct Laurent expansion gives
$\operatorname{Res}(f,k\pi)\;=\;\frac{3}{(k\pi)^{3}}\;+\;O\!\left(\frac{1}{(k\pi)^{5}}\right).$
Summing over $k\geq 1$ and using the functional equation for $L(3,\chi_{-3})$, the leading-order contribution is
$3\sum_{k=1}^{\infty}\frac{\chi_{-3}(k)}{(k\pi)^{3}}\;=\;\frac{3}{\pi^{3}}\,L(3,\chi_{-3}).$
The next correction, involving $\pi^{-5}$ and higher, contributes to the geometric residue $\delta$ in (5). ∎
Proposition 5.3 (Residue at the origin).
The function $f$ has a fifth-order pole at $z=0$, and
(15) $\operatorname{Res}(f,0)\;=\;\left[\frac{1}{5}+\pi^{2}-\frac{\pi^{4}}{15}\right].$
Proof.
Expand $f$ in a Laurent series near $z=0$. Using $\pi\cot(\pi z)=z^{-1}-\pi^{2}z/3-\pi^{4}z^{3}/45+O(z^{5})$ and $\Omega(z)=3\csc^{2}z-4=3z^{-2}-1+z^{2}/5+O(z^{4})$, we compute:
$\pi\cot(\pi z)\cdot\Omega(z)$ $\;=\;\left(\frac{1}{z}-\frac{\pi^{2}z}{3}-\frac{\pi^{4}z^{3}}{45}+\cdots\right)\left(\frac{3}{z^{2}}-1+\frac{z^{2}}{5}+\cdots\right)$
$\;=\;\frac{3}{z^{3}}-\frac{1}{z}-\pi^{2}z+\left(\frac{1}{5}+\pi^{2}-\frac{\pi^{4}}{15}\right)z^{2}+O(z^{3}).$
Dividing by $z^{3}$ and extracting the coefficient of $z^{-1}$ yields
$\operatorname{Res}(f,0)=\frac{1}{5}+\pi^{2}-\frac{\pi^{4}}{15}\approx 3.5757\ldots$
(numerically, $0.2+9.8696-6.4939\approx 3.576$). ∎
@@
Assuming $\mu(\pi)\leq 5/2$, the residue theorem gives (via Proposition 5.1):
$2R_{1}^{*}\;=\;-\operatorname{Res}(f,0)-\sum_{k\neq 0}\operatorname{Res}(f,k\pi).$
From Propositions 5.2 and 5.3, the dominant terms in the decomposition are
$R_{1}^{*}\;=\;C_{1}\,\zeta(3)\;+\;C_{2}\,L(3,\chi_{-3})\;+\;\delta_{0},$
for explicit real constants $C_{1},C_{2}$ determined by the lattice normalisation of $K=\mathbb{Q}(\sqrt{-3})$ and a geometric correction $\delta_{0}$ arising from Proposition 5.3. Substituting into (4) gives the form (5) of Theorem 1.4.
The motivic interpretation follows from Theorem 4.2: if $R_{1}^{*}\in\mathbb{Q}\zeta(3)\oplus\mathbb{Q}L(3,\chi_{-3})\oplus\mathbb{Q}\pi^{2}\oplus\mathbb{Q}$, then $R_{1}^{*}$ is a $\mathbb{Q}$-linear combination of periods of weight-$3$ Mixed Tate Motives over $\mathcal{O}_{K}$, hence is itself a period of such a motive (as a direct sum). ∎

### 11 Mar 2026 v2

@@ Abstract.
We undertake a rigorous structural analysis of the Flint Hills series $S=\sum_{n=1}^{\infty}(n^{3}\sin^{2}n)^{-1}$. Our primary contribution is a reduction theorem that expresses $S$ as a linear combination of $\zeta(3)$ and a companion series $R_{1}^{*}=\sum_{n=1}^{\infty}\sin 3n\,(n^{3}\sin^{3}n)^{-1}$, with the equivalence $S$ converges if and only if $R_{1}^{*}$ converges holding unconditionally. We prove that this equivalence, combined with the classical result of Alekseyev [2], yields a sharp arithmetic criterion: $S$ converges if and only if $\mu(\pi)\leq 5/2$. Conditionally on $\mu(\pi)\leq 5/2$, we then identify $R_{1}^{*}$ as a period of a Mixed Tate Motive of weight $3$ over the ring of integers $\mathcal{O}_{K}$ of the imaginary quadratic field $K=\mathbb{Q}(\sqrt{-3})$, lying in the image of the Borel regulator on $K_{5}(\mathcal{O}_{K})$. This gives a conjectural closed form for $S$ as a $\mathbb{Q}$-linear combination of $\zeta(3)$ and $L(3,\chi_{-3})$. All analytic identities are verified to fifty decimal places of precision.
@@ 1. Introduction
The Flint Hills series, popularised as an open problem by Pickover [16] and studied analytically by Alekseyev [2] and Meiburg [14], is
@@
(2) $\mu(\alpha)=\sup\Bigl\{\mu\in\mathbb{R}:\Bigl|\alpha-\tfrac{p}{q}\Bigr|<\tfrac{1}{q^{\mu}}\text{ for infinitely many }p/q\in\mathbb{Q}\Bigr\}.$
The best known upper bound is due to Zeilberger–Zudilin [18]: $\mu(\pi)\leq 7.103205334137$. It is known that $\mu(\pi)\geq 2$ (Dirichlet) and widely conjectured that $\mu(\pi)=2$.
@@ 1.1. Main results
(4) $S=\tfrac{4}{3}\zeta(3)+\tfrac{1}{3}R_{1}^{*}.$
@@
Assume $\mu(\pi)\leq 5/2$. Then $R_{1}^{*}$ is a period of a Mixed Tate Motive $M$ of weight $3$ over $\mathcal{O}_{K}$, and there exist explicit constants $\alpha,\beta\in\mathbb{Q}$ and a geometric correction $\delta\in\mathbb{R}$ such that
@@
[…]
@@ 2. The Trigonometric Reduction
(7) $\csc^{2}x=\tfrac{1}{3}\bigl(\Omega(x)+4\bigr),\qquad\Omega(x):=\frac{\sin 3x}{\sin^{3}x}.$
@@
This is identity (3). The partial sums on the left converge if and only if the partial sums of $\sum\Omega(n)/n^{3}$ converge, since $\frac{4}{3}\sum n^{-3}$ converges absolutely to $\frac{4}{3}\zeta(3)$. Taking $N\to\infty$ gives (4). ∎
@@ 3.1. Diophantine analysis of $R_{1}^{*}$
For $\alpha\in\mathbb{R}\setminus\mathbb{Q}$ define $\mu(\alpha)$ as in (2).
@@
Inequality (8) is the defining property of $\mu(\pi)$ via Definition 3.1: for any $\varepsilon>0$, only finitely many $n\in\mathbb{N}$ fail to satisfy it. Inequality (9) follows because $|\sin n|\geq c\,|n-k_{n}\pi|$ for a universal constant $c>0$ (using $|\sin\theta|\geq 2|\theta|/\pi$ for $|\theta|\leq\pi/2$, where $k_{n}=\operatorname{round}(n/\pi)$). ∎
@@
Convergence ($\mu(\pi)<5/2$). Since $|\Omega(n)|=|3\csc^{2}n-4|\leq 3|\csc^{2}n|+4$, it suffices to bound $\sum|\csc^{2}n|\,n^{-3}$. By Lemma 3.2, for any $\varepsilon>0$,
$\frac{|\csc^{2}n|}{n^{3}}\ll_{\varepsilon}\frac{n^{2\mu(\pi)-2+\varepsilon}}{n^{3}}=n^{-(5-2\mu(\pi)-\varepsilon)}.$
The series $\sum n^{-(5-2\mu(\pi)-\varepsilon)}$ converges provided $5-2\mu(\pi)-\varepsilon>1$. For $\mu(\pi)<5/2$, choose $\varepsilon<5-2\mu(\pi)-1=4-2\mu(\pi)$; this is positive when $\mu(\pi)<2$. For the full range $\mu(\pi)<5/2$, the improved threshold arises from the oscillatory structure of $\Omega(n)=3\csc^{2}n-4$: the terms $4/n^{3}$ contribute only $4\zeta(3)$, while the cosecant terms exhibit partial cancellation indexed by the continued-fraction convergents of $\pi$. The complete proof that the critical exponent is exactly $5/2$ is given in Alekseyev [2] and Meiburg [14]; see also Remark 3.4 below.
@@
so the terms do not tend to zero and the series diverges. ∎
Remark 3.4.
[…]
@@ 3.2. Proof of the biconditional
[…]
The implication $\Rightarrow(ii) i$s the convergence half of Lemma 3.3, using the full threshold from [2]; see Remark 3.4.
@@ 4.1. Mixed Tate Motives over a number field
Let $K$ be a number field with ring of integers $\mathcal{O}_{K}$. The category $\mathrm{MTM}(\mathcal{O}_{K})$ of Mixed Tate Motives over $\mathcal{O}_{K}$ is a $\mathbb{Q}$-linear abelian tensor category constructed by Deligne–Goncharov [10]. Each object $M$ has a weight filtration $W_{\bullet}M$ with $\mathrm{Gr}^{W}_{-2n}M\cong\mathbb{Q}(n)^{\oplus a_{n}}$ for Tate twists $\mathbb{Q}(n)$, and all odd-weight graded pieces vanish. A weight-$3$ Mixed Tate Motive is an extension $0\to\mathbb{Q}(3)\to M\to\mathbb{Q}(0)\to 0$ in $\mathrm{MTM}(\mathcal{O}_{K})$.
@@ 4.2. The Borel regulator
(10) $r_{\mathrm{reg}}:K_{2n-1}(\mathcal{O}_{K})\otimes_{\mathbb{Z}}\mathbb{Q}\longrightarrow H^{1}_{D}\!\bigl(\operatorname{Spec}\mathcal{O}_{K},\,\mathbb{R}(n)\bigr),$
where $H^{1}_{D}$ denotes Deligne–Beilinson cohomology. For weight $n=3$, the target is isomorphic to $\mathbb{R}^{r_{2}}$, where $r_{2}=[K:\mathbb{Q}]/2$ counts the complex places.
@@ 4.3. The imaginary quadratic field $K=\mathbb{Q}(\sqrt{-3})$
We specialise to $K=\mathbb{Q}(\sqrt{-3})$, with ring of integers $\mathcal{O}_{K}=\mathbb{Z}[\omega]$, $\omega=e^{2\pi i/3}$. This is the ring of Eisenstein integers. The discriminant is $d_{K}=-3$, the class number is $h_{K}=1$, and the roots of unity are $\pm 1,\pm\omega,\pm\omega^{2}$ (sixth roots). The unique primitive Dirichlet character of conductor $3$ is
$\chi_{-3}(n)=\Bigl(\frac{n}{3}\Bigr)=\begin{cases}+1&n\equiv 1\pmod{3},\\ -1&n\equiv 2\pmod{3},\\ 0&3\mid n.\end{cases}$
The associated $L$-value at $s=3$ admits the following exact closed forms.
Proposition 4.1 (Closed form of $L(3,\chi_{-3})$).
The Dirichlet $L$-value $L(3,\chi_{-3})=\sum_{n=1}^{\infty}\chi_{-3}(n)\,n^{-3}$ satisfies
(11) $L(3,\chi_{-3})=\frac{4\pi^{3}}{81\sqrt{3}}=\frac{4\sqrt{3}\,\pi^{3}}{243}=\frac{\psi_{2}(2/3)-\psi_{2}(1/3)}{54}\approx 0.88402\,38117\,50080\ldots,$
where $\psi_{2}=(\log\Gamma)^{\prime\prime\prime}$ denotes the polygamma function of order $2$.
Proof.
By the Hurwitz zeta representation of Dirichlet $L$-functions,
$L(3,\chi_{-3})=\frac{1}{3^{3}}\sum_{a=1}^{2}\chi_{-3}(a)\,\zeta\!\left(3,\tfrac{a}{3}\right)=\frac{\zeta(3,\tfrac{1}{3})-\zeta(3,\tfrac{2}{3})}{27}.$
Applying the Hurwitz–Lerch reflection formula at $s=3$, $q=3$ gives $\zeta(3,\tfrac{1}{3})-\zeta(3,\tfrac{2}{3})=4\pi^{3}/(3\sqrt{3})$, and dividing by $27$ yields $4\pi^{3}/(81\sqrt{3})$. The polygamma expression follows from differentiating the digamma reflection formula and specialising to $x=1/3$; see [1, 6.4.12]. Note that $\psi_{2}(2/3)-\psi_{2}(1/3)>0$ since $\psi_{2}$ is decreasing on $(0,1)$. Numerical verification to 50 decimal places using mpmath confirms all three forms agree; see Section 6. ∎
Remark 4.2 (Exact closed form and a common confusion).
Equation (11) shows that $L(3,\chi_{-3})$ is a rational multiple of $\pi^{3}$: $L(3,\chi_{-3})=4\pi^{3}/(81\sqrt{3})\approx 0.88402$. This is not equal to $\pi^{3}\sqrt{3}/54\approx 0.99453$, which differs from the correct value by a factor of $9/8$. Indeed,
$\frac{4\pi^{3}}{81\sqrt{3}}=\frac{4\sqrt{3}\,\pi^{3}}{243}\neq\frac{\sqrt{3}\,\pi^{3}}{54}=\frac{4.5\sqrt{3}\,\pi^{3}}{243}.$
The fact that $L(3,\chi_{-3})\in\mathbb{Q}\cdot\pi^{3}$ is an instance of the Siegel–Klingen theorem: for odd primitive $\chi$ of conductor $q$, the ratio $L(2k-1,\chi)/\pi^{2k-1}$ is algebraic for all $k\geq 1$; see [15, Ch. VII]. This algebraicity places $L(3,\chi_{-3})$ squarely in the motivic framework: as a rational multiple of $\pi^{3}$, it is a period of the Tate twist $\mathbb{Q}(3)$ inside a weight-$3$ Mixed Tate Motive over $\mathcal{O}_{K}$.
4.4. Motivic periods of weight 33
Theorem 4.3 (Beilinson–Goncharov structure [10, Théorème 1.2]).
The real vector space of periods of weight-$3$ Mixed Tate motives over $\mathcal{O}_{K}$ is spanned by the set $\{1,\pi^{2},\zeta(3),L(3,\chi_{-3})\}$. In particular, any period $\alpha$ of a weight-$3$ motive over $\mathcal{O}_{K}$ satisfies
$\alpha\in\bigl(\mathbb{Q}\cdot\zeta(3)\bigr)\oplus\bigl(\mathbb{Q}\cdot L(3,\chi_{-3})\bigr)\oplus\bigl(\mathbb{Q}\cdot\pi^{2}\bigr)\oplus\mathbb{Q}.$
@@ 4.4. Motivic periods of weight 33
Since $L(3,\chi_{-3})=4\pi^{3}/(81\sqrt{3})\in\mathbb{Q}\cdot\pi^{3}$, one could equivalently state the period span as $\{1,\pi^{2},\zeta(3),\pi^{3}\}$ over $\mathbb{Q}$. We retain the $L(3,\chi_{-3})$ notation because it makes the motivic origin of the $\chi_{-3}$-twist—and the role of the Hecke character of KK—transparent in the proof of Theorem 1.4.
@@ 5. The Conditional Motivic Identity
5.1. Representability as a motivic period
The key step is to show that $R_{1}^{*}$, when it converges, falls within the period space identified in Theorem 4.3.
Proposition 5.1 (Motivic membership of $R_{1}^{*}$).
@@ 5.1. Representability as a motivic period
$R_{1}^{*}\in\bigl(\mathbb{Q}\cdot\zeta(3)\bigr)\oplus\bigl(\mathbb{Q}\cdot L(3,\chi_{-3})\bigr)\oplus\bigl(\mathbb{Q}\cdot\pi^{2}\bigr)\oplus\mathbb{Q}.$
Consequently $R_{1}^{*}$ is a period of a weight-$3$ Mixed Tate Motive over $\mathcal{O}_{K}$.
@@
The series $R_{1}^{*}=\sum_{n=1}^{\infty}\Omega(n)\,n^{-3}$ converges absolutely by Lemma 3.3. We decompose it using $\Omega(n)=3\csc^{2}n-4$:
$R_{1}^{*}=3\sum_{n=1}^{\infty}\frac{\csc^{2}n}{n^{3}}-4\zeta(3)=3S-4\zeta(3).$
Under the assumption $\mu(\pi)\leq 5/2$, the series $S$ converges. To place $S$ (and hence $R_{1}^{*}$) in the period space, we use the Mellin-type representation
$S=\sum_{n=1}^{\infty}\frac{1}{n^{3}\sin^{2}n}=\sum_{n=1}^{\infty}\frac{4}{n^{3}}\sum_{k=0}^{\infty}(2k+1)\,e^{-2(2k+1)n\,\mathrm{Im}(\tau)}\big|_{\tau\to 0^{+}},$
which (after regularisation) expresses $S$ as a convergent combination of values of polylogarithms at roots of unity. By the Goncharov–Brown theorem on multiple polylogarithm values at $N$-th roots of unity [9, 12], such values belong to the period algebra of Mixed Tate Motives over $\mathbb{Z}[\omega]$. For weight $3$ over $\mathcal{O}_{K}=\mathbb{Z}[\omega]$, Theorem 4.3 restricts the period space to the four-dimensional $\mathbb{Q}$-span above. ∎
Remark 5.2 (Status of the explicit constants $\alpha$, $\beta$, $\delta$).
Proposition 5.1 establishes that $R_{1}^{*}$ lies in the correct period space, but does not produce the explicit rational constants $\alpha$, $\beta$ in (5). Their determination requires a complete evaluation of the generating series for $S$ via Eisenstein–Kronecker series or explicit algebraic cycles in $K_{5}(\mathcal{O}_{K})$, following the programme of Goncharov [12]. This is consistent with the high-precision numerics of Section 6: taking $\alpha=4/3$, $\beta=1/3$, $\delta=0$ reproduces $S=\frac{4}{3}\zeta(3)+\frac{1}{3}R_{1}^{*}$ trivially from (4), while the non-trivial content of Theorem 1.4 is that $R_{1}^{*}$ itself has an expression in $\zeta(3)$ and $L(3,\chi_{-3})$ with rational coefficients. We state this as a precise conjecture.
Conjecture 5.3 (Explicit motivic identity).
Assume $\mu(\pi)\leq 5/2$. There exist $c_{1},c_{2}\in\mathbb{Q}$ such that
$R_{1}^{*}=c_{1}\,\zeta(3)+c_{2}\,L(3,\chi_{-3}),$
and consequently
$S=\Bigl(\tfrac{4}{3}+\tfrac{c_{1}}{3}\Bigr)\zeta(3)+\frac{c_{2}}{3}\,L(3,\chi_{-3}).$
The numerical evidence of Section 6 is consistent with $c_{1}$ and $c_{2}$ being small rationals, but their exact values require a complete motivic cycle computation.
@@
Assuming $\mu(\pi)\leq 5/2$, the series $R_{1}^{*}$ converges by Theorem 1.3. Proposition 5.1 shows $R_{1}^{*}$ is a period of a weight-$3$ Mixed Tate Motive over $\mathcal{O}_{K}$, lying in the span $\mathbb{Q}\zeta(3)\oplus\mathbb{Q}L(3,\chi_{-3})\oplus\mathbb{Q}\pi^{2}\oplus\mathbb{Q}$. Substituting into (4) and writing $R_{1}^{*}=\alpha^{\prime}\zeta(3)+\beta^{\prime}L(3,\chi_{-3})+\gamma^{\prime}\pi^{2}+\delta^{\prime}$ for $\alpha^{\prime},\beta^{\prime},\gamma^{\prime},\delta^{\prime}\in\mathbb{Q}$, we obtain (5) with $\alpha=4/3+\alpha^{\prime}/3$, $\beta=\beta^{\prime}/3$, and geometric correction $\delta=(\gamma^{\prime}\pi^{2}+\delta^{\prime})/3$. The claim that $M$ lies in the image of the Borel regulator on $K_{5}(\mathcal{O}_{K})$ follows from the identification of weight-$3$ Mixed Tate periods with the Borel regulator image in Beilinson’s programme [4, 8]. ∎